Find the formula for he parabola with the given point of (2,36) and the zeros (-4,0) and (5,0)
Answer by stanbon(75887)

Find the formula for he parabola with the
given point of (2,36) and the zeros (-4,0) and (5,0)
----
y = a(x+4)(x-5)
Solve for "a" using (2,36)
36 = a(-2)(-3)
36 = a(6)
a = 6
------
Equation::
y = 6(x+4)(x-5)
=======================
Cheers,
Stan H.